The price S(u), 0 􏰃 u 􏰃 t, of the share is driven by a geometric Brownian motion: S(u) = Seμu+σW(u). A proportional dividend on this share is paid continuously at rate q > 0 and is reinvested in the share. The continuously compounded interest rate is r. Compute the no-arbitrage price of a derivative with expiration time t and payoff function

R(t) = [S(t/3)S(2t/3)S(t)]1/3 .



1
Expert's answer
2021-05-11T13:57:14-0400

Consider the Brownian motion given by


"S(u)=Se^{\u03bcu+\u03c3W(u) }" "0\u2264u\u2264t \u2026\u2026(a)"

Payoff function is given by

"R(t)=[S(\frac{t}{3})S(\frac{2t}{3})S(t)^{\frac{1}{3}} ……(b)"

where t is the expiration time.

For this first find the value of "S \space at \space u=0,t,(\frac{t}{3}),(\frac{2t}{3})"

"At \space u=0"

equation(a) becomes

"S(0)=Se^{0+\u03c3W(0)}"

"S(0)=Se^{\u03c3W(0) } \u2026\u2026(c)"

at

 u=t

"S(t)=Se^{\u03bct+\u03c3W(t) } \u2026\u2026(d)"

"at u=\frac{t}{3}"

"S(\frac{t}{3})=Se^{μ(\frac{t}{3})+σW(\frac{t}{3}) } ……(e)"

"at u=\frac{2t}{3}"

"S(\frac{2t}{3})=Se^{μ(\frac{2t}{3})+σW(\frac{2t}{3}) } ……(f)"

Put the values from equation(c) ,equation(d) ,equation(e) and equation(f)in equation (b)

"R(t)=[S(\frac{t}{3})S(\frac{2t}{3})S(t)]^{\frac{1}{3}}"


"R(t) =[(Se^{μ(\frac{t}{3})+σW(\frac{t}{3})})(Se^{μ(\frac{2t}{3})+σW(\frac{2t}{3})})(Se^{μt+σW(t)})]^\frac{1}{3}"

"=[(S^{3}e^{μ(\frac{t}{3})+σW(\frac{t}{3})})(e^{μ(\frac{2t}{3})+σW(\frac{2t}{3})})(e^{μt+σW(t)})]^\frac{1}{3}"

"=S[(e^{μ(\frac{t}{3})+σW(\frac{t}{3})})(e^{μ(\frac{2t}{3})+σW(\frac{2t}{3})})(e^{μt+σW(t)})]^\frac{1}{3}"

"=S[ e^{μ}(\frac{t}{3}+\frac{2t}{3}+t)eσW(\frac{t}{3}+\frac{2t}{3}+t) ]^{\frac{1}{3}}"

"=S[e^{μ(2t)}e^{σW(2t)}]^\frac{1}{3}"

"R(t)=S[e^{μ(2t)+σW(2t)}]^\frac{1}{3 } ……(g)"


Now we find no-arbitrage price of a derivative .

differentiate equation(g) with respect to t.

"R'(t)=\frac{S}{3}[e^{μ(2t)+σW(2t)}]^{−\frac23}(e^{μ(2t)+σW(2t})(2μ+2σW)"

"=\frac{S}{3}[e^{μ(2t)+σW(2t})]^\frac{1}{3} (2μ+2σW)"


"R'(t)=\frac{2S}{3}[e^{μ(2t)+σW(2t)}]^\frac{1}{3} (μ+σW)"